OpenStudy (anonymous):

According to the Bohr model of hydrogen, the wavelength of the photon emitted when the electron in hydrogen jumps from the outermost circular orbit (n = oo) to the innermost circular orbit (n = 1). Please help I am confused with the Bohr model

OpenStudy (jfraser):

plug into the rydberg equation\[\frac{1}{\lambda} = R(\frac{1}{n_1} - \frac{1}{n_2})\] since \(n_2\) is infinite, \(\frac{1}{n_2}\) is essentially zero

OpenStudy (anonymous):

Thank You